OpenStudy (anonymous):

What is the solution for x in the proportion three over five equals twelve over x

OpenStudy (imstuck):

Here you would cross-multiply. 3 times x = 5 times 12. 3x = 60. Now divide by 3.

OpenStudy (anonymous):

It equals 20

OpenStudy (anonymous):

thank you!!!!

OpenStudy (anonymous):

3x=5*12. 5*12=60. then divide by 3: 60/3=20